# Universal Approximation Theory: Foundations for Parallelism in Neural Networks

### v1

@@ 4.1 Dataset
4.2 The Feasibility of Para-Former
To validate the feasibility and characteristics of Para-Former, we conducted training and validation on several commonly used image classification datasets, with all models trained for 500 epochs. The experimental design involved preprocessing inputs using the common ViT techniques of patch extraction and positional encoding. We then used Transformer as the basic model architecture, setting all network depths to 1 and configuring Para-Former from single-layer to multi-layer setups, as shown in Figure 3.a. A single-layer setup corresponds to a ViT with a depth of 1. The results of the models are presented in Table 2.
It is evident from Table 2 that as the number of parallel network layers increases, the model accuracy improves progressively (Except for the result of Para-F-1-24 being worse than Para-F-1-12 on the Flower-102 dataset, this variation is within a reasonable range.). According to UAT theory, increasing the network layers is equivalent to increasing $N$ in Equation (1), thereby enhancing the approximation capability of UAT. The results in Table 2 are in complete agreement with UAT theory. Therefore, we believe that this parallel network approach is entirely feasible.
However, the overall prediction accuracy in Table 2 is quite low. What could be the reasons for this? We believe there are three main factors affecting the model’s predictive performance:
1. Model’s Fitting Capability: This is influenced by the model’s size and design. Specifically, the value of $N$ in UAT and the degrees of freedom in the parameters. The larger the $N$ value, the stronger the UAT’s fitting capability. Parameter freedom refers to the number of parameters affected by the input in UAT, such as a freedom of 1 in Eq. 8 and a freedom of 6 in Eq. 9. The degrees of freedom correspond to the model’s ability to dynamically fit based on input. Merely increasing $N$ enhances the model’s ability to fit specific functions, but images are often diverse. Therefore, the network must have the ability to dynamically fit the corresponding function based on the input. Higher degrees of freedom mean the model can fit a wide range of features with significant differences. Additionally, the bias term in the network can also dynamically fit based on input using UAT, which we refer to as bias freedom. Models with bias freedom have a stronger dynamic fitting capability. The mathematical model generated by the residual network is an example of UAT with bias freedom, achieving the fitting capability of a multi-layer network with just shallow layers. We will talk about this problem in Section 4.3.
2. Data Influence: Data is another critical factor in deep learning training. Since natural images are usually diverse, limited data can lead to overfitting. For example, in a cat-dog classification task, if dogs in the training set mostly appear against backgrounds of blue skies and grass, these backgrounds might be mistakenly considered part of the dog’s features, a phenomenon known as feature coupling. If cats rarely appear in such backgrounds in the training set but do in the test set, they might be misclassified as dogs. Training on a large, diverse dataset means that common backgrounds frequently appear with various objects or animals, such as cats, dogs, and elephants against blue skies, white clouds, houses, and trees. Thus, even when classifying objects, the model inherently learns background information. In this context, fine-tuning on the cat-dog dataset yields better results because the background does not interfere with dog feature recognition. Without sufficient data, even a network with strong fitting capabilities cannot perform well and is prone to overfitting. This problem will be considered in Section 4.4.
Considering these factors, it is clear that simply improving the model architecture is not enough. A robust training dataset is equally important for achieving high predictive accuracy.
Dataset Para-F-6-6 Para-F-2-24 Para-F-3-24 Para-F-6-24
@@ 4.2 The Feasibility of Para-Former
4.3 Model Depth Effect
Based on the conclusions from Section 4.2, we will explore the impact of network depth on predictive performance in this section. According to Eq. 8 and Eq. 9, we know that model depth primarily affects the degrees of freedom of the model parameters in UAT. Therefore, we designed Para-Former models with 24 layers and varying depths of 2, 3, and 6, and compared their predictive performance with that of a ViT with a depth of 6. The results are shown in Table 3.
It is evident that all results surpass those of Para-F-1-24 in Table 2 because the networks in Table 3 have higher degrees of freedom in their parameters. Notably, Para-F-6-6 consistently outperforms Para-F-2-24 and Para-F-3-24, indicating that the higher the degrees of freedom of the parameters, the stronger the model’s performance. However, the model cannot be too deep, so we must balance network depth and layer count. Another issue is that almost all models hit a performance bottleneck. Particularly, the prediction accuracy for OxfordIIITPet and Flower-102 is very poor. This is because these datasets are relatively small and exhibit high data diversity.
@@ 4.4 Dataset Effect
Table 4: This table shows the results of fine-tuning using Google’s official pre-trained parameters. The pre-trained parameters were obtained from training the ViT model on the ImageNet dataset.
Dataset Cifar-10 Cifar-100 STL-10 OxfordIIITPet Flower-102
@@
Current deep learning models generally perform well on the aforementioned datasets. However, as shown in Table 3, the prediction accuracy of Para-Former is not very high, especially on the OxfordIIITPet and Flower-102 datasets. What is the fundamental reason for this poor performance? In this section, we will explain the relationship between deep learning and data.
First, the results were obtained using the validated ViT model (Para-F-1-1 and Para-Former-6-6), which performed similarly to Para-Former. Given that the ViT model has been extensively validated and is a commonly used model architecture, the effectiveness of the models is unquestionable. So, why is the performance so poor? The root cause lies in the data itself: the model’s performance is generally directly proportional to the amount of available data. Take the Flower-102 dataset, for example, which includes 102 classes of flowers but has only 1,020 images in the training set, with some reserved for validation. This is far from sufficient to train deep learning models. Why, then, does the OxfordIIITPet dataset, which has more data per class on average than Flower-102, perform even worse? This is due to the diversity of the data (i.e., the similarity of features within the same class) and the variety of backgrounds.
Figures 4 and 5 show some results from the OxfordIIITPet and Flower-102 datasets. Both figures indicate that prediction results are worse for categories with diverse backgrounds or inherent diversity, as seen in Figures 4.b and 5.b. Conversely, results are relatively better when the background is consistent but the species vary, as shown in Figures 4.a and 5.a. The best results are achieved when both the background and species variation are minimal, as shown in Figures 4.c and 5.c. Although the Flower-102 dataset has fewer images, its backgrounds are more consistent and there is less variation among the species. In contrast, animals in the OxfordIIITPet dataset can appear in various forms, while flowers in the Flower-102 dataset are usually photographed from relatively fixed angles, leading to better performance.
Does this mean that deep learning models are not powerful enough or lack generalization and reasoning capabilities? On the contrary, achieving around 25% accuracy on OxfordIIITPet and about 40% on Flower-102 with such limited data actually demonstrates the strength and generalization capability of deep learning models. A common misconception about deep learning is that humans, even if unfamiliar with certain species, can accurately distinguish OxfordIIITPet and Flower-102 images by carefully examining them, thus believing that deep learning lacks generalization or human-like inferencing ability. This viewpoint is flawed. Assuming the human brain functions as a UAT model, it can automatically disregard background information before seeing these images, as this background is frequently encountered in daily life, and the UAT in the human brain has already fitted this information. Therefore, humans only need to focus on specific information and then use it to train their brains to find similarities and differences. For instance, a 10-year-old child’s brain, observing the world at 60 frames per second, would have processed about $1\times 10^{10}$ high-resolution images by the age of 10, many of which include common objects and backgrounds in daily life. Therefore, criticizing deep learning models for being unable to train on small datasets and lacking universality and generalization is unfair.
Thus, we should train on large datasets and then fine-tune on smaller ones. This is also the current deep learning training method, showcasing the powerful generalization capability of deep learning models, especially when there is some correlation between datasets. Following this fine-tuning approach, we fine-tuned the pre-trained ViT model on the aforementioned datasets for just 10 epochs. The results, as shown in Table 4, indicate a significant improvement in prediction performance.
Figure 6 shows the rationality behind fine-tuning. Since we essentially adjust parameters based on the input to produce target results using UAT, if there is a correlation in the data, the corresponding parameters derived from the input ($\mathbf{\alpha}_{i},\mathbf{W}_{i}$, and $\mathbf{\theta}_{i};i=1,\cdot,N$) will be similar. Therefore, fine-tuning is feasible.

### v4

@@ 3.2 Para-former
Table 2: This table shows the predictions of Para-former-1-1, Para-former-1-6, Para-former-1-12, Para-former-1-24 on the datasets CIFAR-10, CIFAR-100, STL-10, OxfordIIITPet, and Flower-102. We simply represent them as Para-F-1-1, Para-F-1-6, Para-F-1-12 and Para-F-1-24.
3.3 Speed-up Ratio
We have proposed a network parallelization scheme, and in this section, we will compare the running speed of this parallelized network with that of traditional serial networks. It is important to note that due to differences in hardware and implementation methods, providing an objective comparison through experiments can be challenging. Therefore, we will focus on a theoretical analysis. Based on the network structure shown in Figure 3 and 4, the inference time of Para-Former is only affected by its network depth.
Assume we use a Para-Former with a depth of $M$ and a layer count of $L$, compared to a traditional Transformer network with a depth of $N$ layers, where $L>>N$. In this case, the running speed of the parallel network would be $N/M$ times faster than that of the serial network. This is because the speed of the parallel network is solely related to the depth of the Para-Former. As the serial network depth increases, the acceleration effect becomes more pronounced.
4 Experiments
Our experimental approach is to first verify the feasibility of the parallel network theory and its adherence to the UAT approximation properties (The larger the value of $N$, the more powerful the model becomes, resulting in better classification performance.) using several commonly used public image classification datasets (Section 4.2). We then explore the impact of network depth on the Para-Former performance (Section 4.3), followed by an investigation into the influence of the dataset (Section 4.4). We conducted experiments with various Para-Former models, where the naming convention uses the first number to indicate the depth and the second number to indicate the number of layers. For instance, Para-former-1-12 denotes a model with 1 depth and 12 layers.
4.1 Dataset
In pursuit of a comprehensive understanding of the efficacy and practicality of parallel models, we meticulously selected several widely-acknowledged and representative image classification datasets for an integrated comparative analysis. These datasets encompass a diverse array of visual content and span various levels of difficulty, thereby furnishing us with a fertile ground to probe the performance boundaries of parallel models. Specifically, our study centers on the following pivotal datasets: CIFAR-10, CIFAR-100 (Krizhevsky 2009), STL-10 (Coates, Ng, and Lee 2011), Flowers-102 (Nilsback and Zisserman 2008) and Oxford-IIIT Pets (Parkhi et al. 2012). In Table 1, we provide more details for these datasets.
Para-F-6-6 Para-F-2-24 Para-F-3-24 Para-F-6-24
@@ 4.1 Dataset
4.2 The Feasibility of Para-Former
To validate the feasibility and characteristics of Para-Former, we conducted training and validation on several commonly used image classification datasets, with all models trained for 500 epochs. The experimental design involved preprocessing inputs using the common ViT techniques of patch extraction and positional encoding. We then used Transformer as the basic model architecture, setting all network depths to 1 and configuring Para-Former from single-layer to multi-layer setups, as shown in Figure 3 (We evaluated the models’ performance using the accuracy metric, calculated as $\text{accuracy}=100\times\frac{\text{correct}}{\text{total}}$). A single-layer setup corresponds to a ViT with a depth of 1. The results of the models are presented in Table 2.
It is evident from Table 2 that as the number of parallel network layers increases, the model accuracy improves progressively (Except for the result of Para-F-1-24 being worse than Para-F-1-12 on the Flower-102 dataset, this variation is within a reasonable range.). According to UAT theory, increasing the network layers is equivalent to increasing $N$ in Eq. (1), thereby enhancing the approximation capability of UAT. The results in Table 2 are in complete agreement with UAT theory. Therefore, we believe that this parallel network approach is entirely feasible.
However, the overall prediction accuracy in Table 2 is quite low. What could be the reasons for this? We believe there are two main factors affecting the model’s predictive performance:
1. Model Fitting Capability: This is influenced by the model’s size and design, particularly the value of $N$ in the UAT and the degrees of freedom of the parameters. A larger $N$ enhances the UAT’s ability to fit functions. The degrees of freedom refer to the number of parameters in the UAT that are influenced by the input. For instance, the degrees of freedom in Equation (8) is 2, while in Equation (9), it is 6 (see Appendix B). The degrees of freedom represent the model’s ability to dynamically fit corresponding functions based on the input. Simply increasing $N$ can improve the model’s ability to fit specific functions, but images are usually diverse. Thus, the network must have the capability to dynamically fit different functions based on the input. Higher degrees of freedom allow the model to fit multiple different functions according to the input. Additionally, the degrees of freedom vary within the Transformer: they mainly consist of weights and biases. The degrees of freedom for weights stem from the MHA, primarily putting linear transformations based on the input. The degrees of freedom for biases come from residual connections, which can dynamically fit the bias terms using UAT. The deeper the residual layers, the more UAT layers correspond to the bias terms, leading to stronger dynamic fitting capabilities. As a result, a shallow serial network can achieve the fitting ability of a multi-layer parallel network with a depth of 1. We will explore this issue further in Section 4.3.
2. Data Influence: Data is another crucial factor in deep learning training. Since natural images are typically diverse, limited data can lead to overfitting. For example, in a cat-dog classification task, if dogs in the training set mostly appear against backgrounds of blue skies and grass, these backgrounds might be mistakenly considered features of dogs. If cats rarely appear in such backgrounds in the training set but do in the test set, they might be misclassified as dogs. Training on a large, diverse dataset means common backgrounds frequently appear with various objects or animals, such as cats, dogs, and elephants alongside blue skies, white clouds, houses, and trees. Therefore the model will learn background information unconsciously when doing classification tasks. Then fine-tuning on a smaller cat-dog dataset will yield better results because the background won’t interfere with the recognition of dog features. Without sufficient data, even a network with strong fitting capabilities cannot perform well and is prone to overfitting. We will consider this issue in Section 4.4.
@@ 4.2 The Feasibility of Para-Former
4.3 Model Depth Effect
Based on the conclusions from Section 4.2, this section explores the impact of network depth on prediction performance. According to equations (8) and (9), the depth of the model primarily affects the degrees of freedom of the model parameters within the UAT. Therefore, we designed different versions of the Para-Former model with 24 layers and depths of 2, 3, and 6, comparing their predictive performance against a ViT with a depth of 6. The results are shown in Table 3. The degrees of freedom for Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 are 48, 18, 72, 72, and 72, respectively (see Appendix B).
Why does Para-F-6-6, despite having the lowest degrees of freedom, only perform worse than Para-F-6-24? This is because the degrees of freedom encompass both the freedom of weights and biases. The freedom of weights can be understood as linear transformations derived from inputs, whereas the freedom of biases results from the approximation by multiple layers of UAT, which allows for stronger dynamic fitting capabilities. The deeper the layers, the stronger the approximation ability. Thus, Para-F-6-24 performs the best, followed by Para-F-6-6, as the number of UAT layer of their bias terms is largest (see Appendix B).
In summary, deeper models have stronger approximation capabilities, but there is a limit to how deep a model can go, necessitating a balance between network depth and layer count.
@@ 4.4 Dataset Effect
Cifar-10 Cifar-100 STL-10 OxfordIIITPet Flower-102
@@
Table 4: This table shows the results of fine-tuning using Google’s official pre-trained parameters of ViT. The pre-trained parameters were obtained from training the ViT model on the ImageNet dataset.
@@
After increasing the depth of the Para-Former, the results have improved to some extent. However, as shown in Table 3, almost all models encounter a performance bottleneck, especially on the OxfordIIITPet and Flower-102 datasets. What is the fundamental reason for this poor performance? This is likely because these datasets are relatively small and exhibit high data diversity. In this section, we will solve this problem.
First, the experimental results are based on validated ViT models (Para-F-1-1 and Para-Former-6-6 are equal to ViT with 1 and 6 depth.), whose performance is similar to that of the Para-Former. Given that ViT models are widely validated and are among the commonly used model architectures, their effectiveness is beyond doubt. So why is the performance so poor? The fundamental reason lies in the data itself: the model’s performance is typically directly related to the amount of available data. Take the Flower-102 dataset as an example, which contains 102 classes of flowers but only 1,020 training images, with some reserved for validation. This is far from sufficient for training deep learning models. Why is the performance even worse on the OxfordIIITPet dataset, which has a higher average number of images per class? This is due to the diversity of the data (i.e., the similarity of features within the same class) and the variability of the background.
Figures 5 and 6 show some results from the OxfordIIITPet and Flower-102 datasets. Both figures indicate that for categories with diverse backgrounds and inherent feature diversity, the prediction results are poor, as shown in Figures 5.b and 6.b. In contrast, when the background is consistent but the object morphology is diverse, the results are relatively better, as shown in Figures 5.a and 6.a. The best results are achieved when both the background and object morphology are minimal, as shown in Figures 5.c and 6.c. Although the Flower-102 dataset has fewer images, its backgrounds are more consistent, and the differences between species are smaller. Conversely, animals in the OxfordIIITPet dataset can appear in various forms, whereas flowers in the Flower-102 dataset are usually photographed from relatively fixed angles, leading to better performance.
Does this mean that deep learning models are not powerful enough or lack generalization and reasoning capabilities? Quite the opposite. With such limited data, deep learning models achieved about 25% accuracy on OxfordIIITPet and about 40% on Flower-102, actually demonstrating the strength and generalization ability of deep learning models. A common misconception about deep learning is that even if people are unfamiliar with certain species, by careful examination, they can accurately distinguish between OxfordIIITPet and Flower-102 images, leading to the belief that deep learning lacks generalization or human-like reasoning capabilities. This view is flawed. Suppose the human brain operates like a UAT model; it can automatically ignore background information before seeing these images because such backgrounds are frequently encountered in daily life, and the UAT in the brain has already fitted this information. Therefore, people only need to focus on specific information and then use this information to train the brain to identify similarities and differences. For example, a 10-year-old child’s brain, observing the world at 10 frames per second (Read and Meyer 2000), would have processed over $1.5\times 10^{9}$ high-resolution images by age 10 (Lu et al. 2017), many containing common objects and backgrounds from daily life. Hence, it is unfair to criticize deep learning models for being unable to train on small datasets and for lacking universality and generalization ability.
Therefore, we should train on large datasets and then fine-tune on smaller ones. This is also the current method for deep learning training, demonstrating the powerful generalization capabilities of deep learning models, especially when there is some correlation between the datasets. Following this approach, we fine-tuned a pre-trained ViT model using only 10 epochs. The results, as shown in Table 4, demonstrate significant improvements in prediction performance.
Figure 7 illustrates the rationale for fine-tuning. The goal, whether in training or prediction, is to obtain the target results based on the input. This can be understood as function fitting, expressed as $\mathbf{y}=\mathbf{G}(\mathbf{x}|\mathbf{A},\mathbf{W},\mathbf{\Theta})$, where $\mathbf{A}={\mathbf{\alpha}_{1}...\mathbf{\alpha}_{N}}$, $\mathbf{W}={\mathbf{W}_{1}...\mathbf{W}_{N}}$, and $\mathbf{\Theta}={\mathbf{\theta}_{1}...\mathbf{\theta}_{N}}$. The purpose of training is to determine $\mathbf{A},\mathbf{W},\mathbf{\Theta}$, which can dynamically change based on the input.
This means that if the background is similar but the objects are different, the resulting parameters will differ. If the model can make correct predictions, it implies that the model can distinguish between the background and the animals. If it could not be distinguished, it might incorrectly use the background as a feature of the animals, leading to the same classification. Moreover, if the model can correctly classify the animals, it indicates that the model can identify the features of the animals.
When the backgrounds are different, because natural backgrounds are generally similar, extensive training can differentiate background information. Thus, if the model can make right predictions, it suggests that the model has comprehensively grasped the animal features, even with significant environmental variations. Training on ImageNet follows this process. A new dataset is just a new series of combinations of various similar backgrounds and animals which have been seen in ImageNet. If the model has already learned sufficient background and animal information on ImageNet, it can dynamically fit the results based on the input. Therefore, only minor adjustments specific to new data are required during fine-tuning.